# Input Reduction Enhanced LLM-based Program Repair

### v1

@@ Abstract.
For targeted evaluation, we constructed LFTBench, the first long-input APR benchmark with 200 real bugs from 20 programming tasks, each paired with a failure‑inducing input whose median size is 1 MB. On this benchmark, ReduceFix shrinks inputs by 89.1% on average and improves overall pass@10 by up to 53.8% relative to a prompt that includes the original test, and by 17.6% compared with omitting the test entirely. Adding the same reduction step to ChatRepair increases its fix rate by 21.3% without other changes. Ablation studies further highlight the impact of input length and compressed failure information on repair success. These results underscore that automatically reducing failing inputs is a practical and powerful complement to LLM-based APR, significantly improving its scalability and effectiveness.
copyright: none
@@ 1. Introduction
APR aims to automatically generate bug-fixing patches for software defects, thereby reducing the manual effort required for debugging (Yang et al., 2025a; Le Goues et al., 2019; Xu et al., 2024). Recently, LLMs have been applied to APR with promising results (Xia and Zhang, 2022; Luo et al., 2025). Many recent APR systems enhance patch generation by including test inputs in the prompt, among which the test suite plays a particularly important role (Kong et al., 2024; Tang et al., 2024; Yang et al., 2024). It provides a concrete example of how the program succeeds or fails, helping the LLM focus on the underlying issue and produce a correct fix. This strategy has shown strong results in recent systems such as ChatRepair (Xia and Zhang, 2024). Existing APR studies typically evaluate on benchmarks such as Defects4J (Just et al., 2014) and HumanEval-Java (Jiang et al., 2023), where test inputs are short and rarely exceed a few hundred characters. However, when the test input becomes too long, it is difficult to pinpoint the root cause of the error, known as the “lost-in-the-middle” phenomenon (Liu et al., 2023; Pham et al., 2015), where information buried in a long prompt receives little attention and overall task performance drops (Yang et al., 2024; Tian et al., 2023). Therefore, an automatic test input reduction step becomes essential before repair.
@@
Syntax-based approaches, such as HDD (Misherghi and Su, 2006) and Perses (Sun et al., 2018), are built on the classical ddmin algorithm (Zeller and Hildebrandt, 2002) but incorporate grammar or tree structure to ensure syntactic validity during reduction. However, adapting these APR tools to new tasks requires designing a new grammar and tuning heuristics for each input format, which is both time-consuming and error-prone. Other techniques, such as ddSMT (Niemetz and Biere, 2013) and J-Reduce (Kalhauge and Palsberg, 2021), target specific domains, but still require domain knowledge and significant manual effort to implement and are not reusable across different formats. In LLM-based APR, the input format varies widely across tasks, often involving different formats ranging from plain text to structured JSON or domain-specific encoding (Liu et al., 2020). As a result, relying on hand-crafted reducers is not scalable and makes automation difficult.
@@
  • Lack of length-aware handling for test input. Although many APR systems embed the full test input into the prompt, few consider how input length affects patch quality. CREF (Yang et al., 2024) finds that on Bard, prompting with the failing test can hurt repair success, achieving 10% lower accuracy than the no-test baseline in some cases, primarily because long inputs overwhelm the LLM and trigger the “lost-in-the-middle” effect (Liu et al., 2023). However, existing APR systems treat the test input as a fixed block of context and have not attempted to shorten or distill it before repair.
  • Limitation of manually crafted input reducers. Prior approaches on input reduction often relies on handwritten syntax grammars or domain-specific rules, which require significant human effort and deep domain knowledge (Misherghi and Su, 2006; Sun et al., 2018; Niemetz and Biere, 2013). Even after laborious manual effort, each reducer remains tightly coupled to a specific task or file structure and cannot be generalized across formats. As LLM-based APR must handle a wide range of input types, including plain text, JSON, and custom encoding, manual reducers do not scale, and no existing method supports automatic reducer generation across diverse tasks.
@@
To enable rigorous and leakage‑free evaluation, we build LFTBench, the first APR benchmark that focuses on long test inputs. It contains 200 buggy codes from 20 AtCoder tasks released after the training cut‑off dates of the evaluated LLMs. On LFTBench, ReduceFix with Qwen‑Plus generates syntactically correct reducers for all the 200 bugs, and 95.0% of those reducers successfully shrink the failure‑inducing input by an average of 89.1%. ReduceFix with 4 selected LLMs increases their overall pass@10 by up to 53.8% compared with prompts that embed the whole test. Replacing the reduced input for the original input in ChatRepair (Xia and Zhang, 2024) raises its pass@10 by 21.3%, confirming that ReduceFix can be plugged into existing APR pipelines for an immediate accuracy boost.
@@
  • Benchmark. We release the first APR benchmark, LFTBench, which contains 200 bugs across 20 different real-world programming tasks, including long test inputs.
  • Comprehensive evaluations. Extensive experiments on LFTBench demonstrate that ReduceFix is able to generate reducers for each task, with 95.0% of those reducers successfully compressing the corresponding failure-inducing input, achieving an average size reduction of 89.1%. Across 4 LLMs with different sizes, the reduced test inputs increase pass@10 by up to 53.8% relative to both the baseline and the original tests. We also conduct an ablation study that varies prompt length and the presence of key test input/output lines. Results show that shortening the test input yields the largest gain, whereas including only an output diff offers little benefit. We further evaluate ReduceFix by integrating it into ChatRepair (Xia and Zhang, 2024), with results indicating that it provides complementary benefits to other APR approaches.
@@ 2. Motivating Example
A single extensive test input can overwhelm an LLM with tokens and conceal the actual defect, which may be located deep within the prompt. This “lost‑in‑the‑middle” effect lowers the attention paid to the critical lines and prevents the model from producing an accurate patch. For example, Problem C of AtCoder Beginner Contest 376111https://atcoder.jp/contests/abc376/tasks/abc376_c makes the issue concrete. The task receives two sorted sequences, $A$ and $B$, and must print the maximum element of $A$ that is not matched by any element of $B$, or $-1$ when more than one element remains unmatched. Figure 1 sketches the whole scenario.
Listing 1 displays the key part of a wrong‑answer submission (No. 65060141). This submission walks the two sequences from the back, increments cnt, and records mx whenever $a[l]$ is larger than $b[r]$. If $|A|>|B|$, the while‑loop exits before the tail of $A$ is checked, so some mismatches slip through.
1while (l >= 1 && r >= 1) {
2 if (b[r] >= a[l]) { r–; l–; }
3 else { cnt++; mx = max(mx, a[l]); l–; }
4}
5if (cnt > 1) { cout << -1; return 0; }
6cout << mx;
Listing 1: Excerpt of the wrong submission.
The online judge provides a failure-inducing input that exceeds 3 MB. When Qwen2.5‑Coder‑7B‑Instruct is prompted with the task statement, the buggy code, and this complete file, it adds only a superficial guard and still ignores the trailing elements, so the program continues to fail. The tokens that expose the oversight sit near the midpoint of the 3 MB prompt, far from either end, where the model focuses most of its capacity.
As shown in Figure 1, ReduceFix addresses this by inserting an automatic reduction stage before repair. It prompts Qwen‑Plus once to write a task‑specific Python script that leverages the classical ddmin algorithm (Zeller and Hildebrandt, 2002) and needs no extra manual effort. Running the LLM-generated reducer reduces the 3 MB test to a three-line counterexample that still forces the buggy and reference programs to diverge.
@@
Long failure inputs therefore hide the defect and mislead the repair model. Applying a reduction restores focus by keeping only the few tokens that matter, and an LLM can generate the reducer automatically, so the entire process is hands‑free. In this example, the reduced prompt improves repair accuracy from an incorrect patch to a fully accepted solution, showing how length control, systematic reduction, and LLM‑generated tooling work together to overcome the lost‑in‑the‑middle barrier within APR scenarios.
@@ 3.1. Overview
ReduceFix receives five inputs: the task description $P$, a correct reference solution $A$, a buggy submission $s_{w}$, the hidden test suite $I$, and one failure‑inducing input $i_{0}$. Its goal is to shrink the failure‑inducing input $i_{0}$ that still distinguishes $A$ from $s_{w}$, then guide an LLM to repair the bug.
The pipeline proceeds in three stages, shown in Figure 2. Reducer Generation prompts a code LLM once and returns a customized reducer script that is enable to automatically reduce the given failing-inducing input for the task. Input Reduction executes the generated reducer script under a time limit to shrink the failing-inducing input $i_{0}$ into a reduced test input $i^{*}$. Patch Generation embeds $\langle P,s_{w},i^{*}\rangle$ in a repair prompt, samples candidate patches, and validates each one against the entire test suite $I$ until a correct program $\hat{s}$ is found or the attempt stops.
@@
3:  $R\leftarrow\text{LLM\_Call}(\Pi)$
4:  if StaticCheckFail$(R)$ then
@@
8:  $i^{*}\leftarrow R.\texttt{reduce}(i_{0})$
9:  if $A(i^{*})=s_{w}(i^{*})$ or Timeout/RE then
10:     return failure
11:  end if
12:  Phase 3: Patch Generation
13:  $\hat{s}\leftarrow\text{LLM\_Call}(P,\;s_{w},\;i^{*})$
14:  if $\forall\,i\in I:A(i)=\hat{s}(i)$ then
15:     return $\hat{s}$
16:  end if
17:  return failure
@@ 3.2. Reducer Generation
Instead of directly reducing test inputs using LLMs, ReduceFix leverages one-shot learning to inject the knowledge of the ddmin algorithm (Zeller and Hildebrandt, 2002) into the LLM. This enables the model to adapt a well-designed reduction algorithm to various input formats and produce effective, customized reducers.
More specifically, ReduceFix builds one prompt that joins the full task statement $P$, a single‑shot example drawn from task ABC330D (Inc., 2025) together with its working reducer, and a few I/O pairs from the current task. The prompt, as shown in Listing 2, is sent to Qwen-Plus (Hui et al., 2024) with a temperature of 00 (greedy decoding). The model returns a reducer $R$, which is a customized reducer derived from the ddmin algorithm (Zeller and Hildebrandt, 2002).
@@
Listing 2: One-shot reducer prompt fed to Qwen-Plus
@@ 3.3. Input Reduction
In this stage, the LLM‑generated reducer will iteratively shrink the given failure-inducing input $i_{0}$ while preserving the failure, i.e., the output inconsistency between the accepted reference solution $A$ and the buggy submission $s_{w}$. More specifically, the reducer begins by dividing the input into chunks and systematically attempts to produce smaller inputs by removing parts of the original input. For a produced smaller input $i$, the reducer checks whether it could still cause the inconsistency between the reference correct code $A$ and buggy code $s_{w}$, i.e., feeding that input $i$ to both $A$ and $s_{w}$ and evaluating whether their outputs are different. If it does, it continues the reduction iteration by replacing the original input with a smaller input; if not, it increases the granularity by splitting the original input into more chunks. This process repeats until no further reduction is possible.
@@ 3.4. LLM-Guided Repair
In this stage, ReduceFix first concatenates the task description $P$, the buggy submission $s_{w}$ and the reduced failing-inducing input $i^{*}$ in a repair prompt, utilizes LLM to samples candidate patches and validates each one until a correct patched program $\hat{s}$ is found.
Because of LLMs’ context length constraints, input prompts have a limited budget. Due to that the reduced test input $i^{*}$ is relatively small, most of them can be inserted into the repair prompt without modification. When $i^{*}$ (or its associated output) still exceeds a configurable budget $L$ lines, ReduceFix truncates the literal text shown to the LLM: it keeps the first $\lceil L/2\rceil$ lines and the last $\lfloor L/2\rfloor$ lines, and never splits a line in the middle. Truncation affects only the prompt; the complete file is preserved for compilation and test execution, so semantic correctness is not compromised.
During candidate patch sampling, ReduceFix utilizes the LLM to generate at most $N=10$ candidate patches. Each candidate must compile within ten seconds and is executed against the full hidden suite $I$. A timeout, runtime error, or wrong answer counts as a failed attempt. The repair step succeeds when the first the patched program $\hat{s}$ passes the entire test suite $I$:
@@
Listing 3 shows the exact prompt template. If truncation occurred, the ellipsis token appears inside the fenced block to signal omitted lines. No other explanatory text is added, keeping the prompt well below typical context limits even on compact LLMs.
1### Problem Description
2{problem_description}
3### Your Incorrect Code
4“‘cpp
5{wa_code}
6“‘
7### Failing Case
8Input:
9“‘
10{reduced_failing_input}
11“‘
12Your Output:
13“‘
14{wa_output}
15“‘
16Expected Output:
17“‘
18{expected_output}
19“‘
20### Your Task
21Fix the C++ code to pass ALL test cases (including hidden ones).
22### Critical Guidelines
231. Focus on algorithmic correctness - NO hard-coded values
[…]
253. Handle edge cases (empty input, single element, max constraints)
264. Use standard C++20 and avoid non-portable extensions
27### Output Format
28Provide ONLY the complete fixed C++ program inside a single cpp block.
Listing 3: LLM repair prompt used in ReduceFix

### v2

@@ Abstract.
For targeted evaluation, we constructed LFTBench, the first long-input APR benchmark with 200 real bugs from 20 programming tasks, each paired with a failure-inducing input whose median size is 1 MB. On this benchmark, ReduceFix shrinks inputs by 89.1% on average and improves overall pass@10 by up to 53.8% relative to a prompt that includes the original test, and by 17.6% compared with omitting the test entirely. Adding the same reduction step to ChatRepair and CREF increases their fix rate by 21.3% and 2.6%, respectively, without other changes. Our gains hold against an ddmin-only reducing template baseline and transfer to repository-level OSS-Fuzz cases. Ablation studies further highlight the impact of input length and compressed failure information on repair success. These results underscore that automatically reducing failing inputs is a practical and powerful complement to LLM-based APR, significantly improving its scalability and effectiveness.
copyright: noneconference: Proceedings of the 48th International Conference on Software Engineering; April 2026; Rio de Janeiro, Braziljournalyear: 2026
@@ 1. Introduction
APR aims to automatically generate bug-fixing patches for software defects, thereby reducing the manual effort required for debugging (yang2025surveyllmbasedautomatedprogram; le2019automated; xu2024aligning; chen2025prometheus). Recently, LLMs have been applied to APR with promising results (xia2022less; luo2025unlocking). Many recent APR systems enhance patch generation by including test inputs in the prompt, among which the test suite plays a particularly important role (kong2024contrastrepair; tang2024code; yang2024cref). It provides a concrete example of how the program succeeds or fails, helping the LLM focus on the underlying issue and produce a correct fix. This strategy has shown strong results in recent systems such as ChatRepair (xia2024automated). Existing APR studies typically evaluate on benchmarks such as Defects4J (just2014defects4j) and HumanEval-Java (jiang_impact_2023), where test inputs are short and rarely exceed a few hundred characters. However, when the test input becomes too long, it is difficult to pinpoint the root cause of the error, known as the “lost-in-the-middle” phenomenon (liu2023lost; pham2015hercules), where information buried in a long prompt receives little attention and overall task performance drops (yang2024cref; tian2023chatgpt). Therefore, an automatic test input reduction step becomes essential before repair.
@@
Syntax-based approaches, such as HDD (HDD) and Perses (Perses), are built on the classical ddmin algorithm (DD) but incorporate grammar or tree structure to ensure syntactic validity during reduction. However, adapting these APR tools to new tasks requires designing a new grammar and tuning heuristics for each input format, which is both time-consuming and error-prone. Other techniques, such as ddSMT (niemetz2013ddsmt) and J-Reduce (DBLP:conf/pldi/KalhaugeP21), target specific domains, but still require domain knowledge and significant manual effort to implement and are not reusable across different formats. In LLM-based APR, the input format varies widely across tasks, often involving different formats ranging from plain text to structured JSON or domain-specific encoding (liu2020efficiency). As a result, relying on hand-crafted reducers is not scalable and makes automation difficult.
@@
  • Lack of length-aware handling for test input. Although many APR systems embed the full test input into the prompt, few consider how input length affects patch quality. CREF (yang2024cref) finds that on Bard, prompting with the failing test can hurt repair success, achieving 10% lower accuracy than the no-test baseline in some cases, primarily because long inputs overwhelm the LLM and trigger the “lost-in-the-middle” effect (liu2023lost). However, existing APR systems treat the test input as a fixed block of context and have not attempted to shorten or distill it before repair.
  • Limitation of manually crafted input reducers. Prior approaches on input reduction often relies on handwritten syntax grammars or domain-specific rules, which require significant human effort and deep domain knowledge (HDD; Perses; niemetz2013ddsmt). Even after laborious manual effort, each reducer remains tightly coupled to a specific task or file structure and cannot be generalized across formats. As LLM-based APR must handle a wide range of input types, including plain text, JSON, and custom encoding, manual reducers do not scale, and no existing method supports automatic reducer generation across diverse tasks.
@@
To enable rigorous and leakage-free evaluation, we build LFTBench, the first APR benchmark that focuses on long test inputs, and we also provide a Python mirror, LFTBench-Py, with wrong-answer Python submissions. LFTBench contains 200 buggy codes from 20 AtCoder tasks released after the training cut-off dates of the evaluated LLMs. On LFTBench, ReduceFix with Qwen2.5-Plus generates syntactically correct reducers for all the 200 bugs, and 95.0% of those reducers successfully shrink the failure-inducing input by an average of 89.1%. ReduceFix with 4 selected LLMs increases their overall pass@10 by up to 53.8% compared with prompts that embed the whole test, and the advantage holds on LFTBench-Py. Beyond AtCoder, we validate on 12 OSS-Fuzz crash instances from five projects. Under Docker-grounded validation with Qwen2.5-Plus, reduced inputs raise the micro-average pass@10 to 41.7% from 25.0% over the no-test Baseline, and from 16.7% over the unreduced Origin Test. Replacing the reduced input for the original input in ChatRepair (xia2024automated) raises its pass@10 by 21.3%, and integrating the same plug-in reducer into CREF increases its pass@10 by 2.6% relative, confirming that ReduceFix can be plugged into existing APR pipelines for an immediate accuracy boost.
@@
  • Benchmark. We release the first APR benchmark, LFTBench, which contains 200 bugs across 20 different real-world programming tasks, and a Python mirror LFTBench-Py for cross-language validation.
  • Comprehensive evaluations. On LFTBench, ReduceFix produces reducers for all bugs and successfully reduces inputs in 95.0% of cases with an average size reduction of 89.1%. Reduced tests raise pass@10 by up to 53.8% across four LLMs. Comparisons to ddmin-only and pure LLM reducers show clear advantages in both success rate and cost. On LFTBench-Py, the reduced tests also outperform both the no-test and full-test prompts. On OSS-Fuzz, reduced inputs improve crash preservation and increase pass@10 from Origin Test’s 16.7% to 41.7% under Docker-grounded validation. We further run ablations on prompt length and evidence composition, finding that shortening the failing input yields the largest gains, while appending only output diffs offers limited benefit.
  • Plug-in integration into existing APR systems. We add ReduceFix as a drop-in reducer to ChatRepair and CREF without changing their logic. ChatRepair’s pass@10 rises by 21.3% and CREF’s pass@10 increases by 2.6% relative, showing that ReduceFix is immediately useful as a portable component.
Figure 2. Overview of ReduceFix.
@@ 2. Motivating Example
A single extensive test input can overwhelm an LLM with tokens and conceal the actual defect, which may be located deep within the prompt. This “lost-in-the-middle” effect lowers the attention paid to the critical lines and prevents the model from producing an accurate patch. For example, Problem C of AtCoder Beginner Contest 376 makes the issue concrete. The task receives two sorted sequences, $A$ and $B$, and must print the maximum element of $A$ that is not matched by any element of $B$, or 1-1 when more than one element remains unmatched. Figure 1 sketches the whole scenario.
The failed submission walks the two sequences from the back, increments cnt, and records mx whenever $a[l]$ is larger than $b[r]$. If $|A|>|B|$, the while-loop exits before the tail of $A$ is checked, so some mismatches slip through.
The online judge provides a failure-inducing input that exceeds 3 MB. When Qwen2.5-Coder-7B-Instruct is prompted with the task statement, the buggy code, and this complete file, it adds only a superficial guard and still ignores the trailing elements, so the program continues to fail. The tokens that expose the oversight sit near the midpoint of the 3 MB prompt, far from either end, where the model focuses most of its capacity.
As shown in Figure 1, ReduceFix addresses this by inserting an automatic reduction stage before repair. It prompts Qwen2.5-Plus once to write a task-specific Python script that leverages the classical ddmin algorithm (DD) and needs no extra manual effort. Running the LLM-generated reducer reduces the 3 MB test to a three-line counterexample that still forces the buggy and reference programs to diverge.
@@
Long failure inputs therefore hide the defect and mislead the repair model. Applying a reduction restores focus by keeping only the few tokens that matter, and an LLM can generate the reducer automatically, so the entire process is hands-free. In this example, the reduced prompt improves repair accuracy from an incorrect patch to a fully accepted solution, showing how length control, systematic reduction, and LLM-generated tooling work together to overcome the lost-in-the-middle barrier within APR scenarios.
@@ 3.1. Overview
ReduceFix receives five inputs: the task description $P$, a correct reference solution $A$, a buggy submission $s_{w}$, the hidden test suite $I$, and one failure-inducing input $i_{0}$. Its goal is to shrink the failure-inducing input $i_{0}$ that still distinguishes $A$ from $s_{w}$, then guide an LLM to repair the bug.
The pipeline proceeds in three stages, shown in Figure 2. Reducer Generation prompts a code LLM once and returns a customized reducer script that is able to automatically reduce the given failure-inducing input for the task. Input Reduction executes the generated reducer script under a time limit to shrink the failure-inducing input $i_{0}$ into a reduced test input $i^{*}$. Patch Generation embeds $\langle P,s_{w},i^{*}\rangle$ in a repair prompt, samples candidate patches, and validates each one against the entire test suite $I$ until a correct program $\hat{s}$ is found or the attempt stops.
@@
[…]
4:if StaticCheckFail(R)(R) then
@@
[…]
9:if Timeout/RE or $A(i^{\dagger})=s_{w}(i^{\dagger})$ or $|i^{\dagger}|\geq|i_{0}|$ then
10:  $i^{*}\leftarrow i_{0}$ {forward the original failing input}
11:else
12:  $i^{*}\leftarrow i^{\dagger}$ {use the reduced input}
13:end if
14:Phase 3: Patch Generation
[…]
16:if $\forall\,i\in I:A(i)=\hat{s}(i)$ then
17:  return $\hat{s}$
18:end if
19:return failure
@@ 3.2. Reducer Generation
Instead of directly reducing test inputs using LLMs, ReduceFix leverages one-shot learning to inject the knowledge of the ddmin algorithm (DD) into the LLM. This enables the model to adapt a well-designed reduction algorithm to various input formats and produce effective, customized reducers.
More specifically, ReduceFix builds one prompt that joins the full task statement $P$, a single-shot example drawn from task ABC330D (ABC330D) together with its working reducer, and a few I/O pairs from the current task. The prompt, as shown in Listing 1, is sent to Qwen2.5-Plus (hui2024qwen2) with a temperature of 0 (greedy decoding). The model returns a reducer $R$, which is a customized reducer derived from the ddmin algorithm (DD).
@@
Listing 1: One-shot reducer prompt fed to Qwen2.5-Plus
@@ 3.3. Input Reduction
In this stage, the LLM-generated reducer will iteratively shrink the given failure-inducing input $i_{0}$ while preserving the failure, i.e., the output inconsistency between the accepted reference solution $A$ and the buggy submission $s_{w}$. More specifically, the reducer begins by dividing the input into chunks and systematically attempts to produce smaller inputs by removing parts of the original input. For a produced smaller input $i$, the reducer checks whether it could still cause the inconsistency between the reference correct code $A$ and buggy code $s_{w}$, i.e., feeding that input $i$ to both $A$ and $s_{w}$ and evaluating whether their outputs are different. If it does, it continues the reduction iteration by replacing the original input with a smaller input; if not, it increases the granularity by splitting the original input into more chunks. This process repeats until no further reduction is possible.
@@ 3.4. LLM-Guided Repair
In this stage, ReduceFix first concatenates the task description $P$, the buggy submission $s_{w}$ and the reduced failure-inducing input $i^{*}$ in a repair prompt, utilizes LLM to samples candidate patches and validates each one until a correct patched program $\hat{s}$ is found.
Because of LLMs’ context length constraints, input prompts have a limited budget. Due to the fact that reduced test input $i^{*}$ is relatively small, most of them can be inserted into the repair prompt without modification. When $i^{*}$ (or its associated output) still exceeds a configurable budget $L$ lines, ReduceFix truncates the literal text shown to the LLM: it keeps the first $\lceil L/2\rceil$ lines and the last $\lfloor L/2\rfloor$ lines, and never splits a line in the middle. Truncation affects only the prompt; the complete file is preserved for compilation and test execution, so semantic correctness is not compromised.
During candidate patch sampling, ReduceFix utilizes the LLM to generate at most $N=10$ candidate patches. Each candidate must compile within ten seconds and is executed against the full hidden suite $I$. A timeout, runtime error, or wrong answer counts as a failed attempt. The repair step succeeds when the first patched program $\hat{s}$ passes the entire test suite $I$:
@@
If truncation occurred in the inferences of patches, the ellipsis token appears inside the fenced block to signal omitted lines. No other explanatory text is added, keeping the prompt well below typical context limits even on compact LLMs.